### v1

# Model-free Legibility: Enhancing Human-Robot Interactions through Implicit Communication and Influence Modulation

Abstract
Communication is essential for successful interaction. In human-robot interaction, implicit communication enhances robots’ understanding of human needs, emotions, and intentions. This paper introduces a method to foster implicit communication in HRI without explicitly modeling human intentions or relying on pre-existing knowledge. Leveraging Transfer Entropy, we modulate influence between agents in social interactions in scenarios involving either collaboration or competition. By integrating influence into agents’ rewards within a partially observable Markov decision process, we demonstrate that boosting influence enhances collaboration or competition performance, while resisting influence diminishes performance. Our findings are validated through simulations and real-world experiments with human participants.
Keywords: Human-Robot Interaction, Implicit communication, Influence modulation
1 Introduction
It is widely recognised that communication is key to successful interaction. Humans communicate with each other through both explicit (direct, deliberate communication over an established channel with clear intent to reach a defined recipient [1]) and implicit channels. Implicit communication is a subtle, indirect mode of conveying information, often relying on context, nonverbal cues, and shared understanding between communicators to convey meaning without explicit verbalization [2]. Implicit communication is particularly crucial for human-robot interaction as it enhances a robot’s ability to proactively understand and respond to human needs, emotions, and intentions, thereby facilitating more natural and effective communication and collaboration between humans and robots. Unfortunately, most current human-robot interaction (HRI) studies focusing on implicit communication explicitly model the intention of human participants [3] [4], or rely on existing intention knowledge [5]. This information is challenging to obtain in general settings.
This paper proposes a method to facilitate implicit communication without the need to explicitly model human participants or rely on pre-existing knowledge. Our approach conceptualizes communication as the degree of influence agents have on one another, employing information-theoretic techniques. Specifically, we use Transfer Entropy (TE) [6] to modulate influence in social interactions. We illustrate our method using a social navigation setting [7], targeting scenarios where the objectives of individual agents may result in a need for either collaboration or competition. In particular, we demonstrate the proposed method in a corridor dilemma setting, a popular scenario in HRI [8] [9]. For our experiments, we define this as a scenario where individuals encounter each other in a narrow passage and where a given agent may wish to meet or avoid another agent. This scenario involves implicit information exchange, where participants attempt to discern each other’s intentions without verbal communication. By observing and reacting to subtle cues such as changes in trajectories, participants estimate each other’s objectives to inform their own actions. Interactions of this form can be complex, especially when both participants act simultaneously, leading to a state of ”resonance” that dissipates as mutual understanding is achieved.
We consider a partially observable Markov decision process (POMDP) setting and augment the rewards of agents to include the TE that boosts or resists influence between agents. We find that boosting influence enhances the other participant’s ability to collaborate or compete with the ego-agent, while resisting influence mainly diminishes their ability to collaborate. These findings hold both in self-play [10] simulations and in real-world experiments with human participants.
2 Related Work
Implicit communication has attracted significant attention in the field of robotics. Che et al. [3] present a planning framework that integrates implicit and explicit communication during mobile robot navigation, using a human behavior model and inverse reinforcement learning to generate communicative actions that enhance robot transparency and efficiency. Li et al. [11] [12] introduce a novel interaction framework that enables communication between humans and assistive devices, leveraging gaze movements to infer human intent. Liang et al. [13] investigate the role of implicature, a form of implicit communication, in HRI using the cooperative card game Hanabi, to demonstrate that human players paired with an implicature AI are significantly more likely to perceive their partner as human. These implicit communication studies typically involve explicit modeling of other participants or rely on pre-existing social knowledge [5]. In contrast, our approach is model-free, and does not require pre-exisiting social knowledge.
Implicit communication is closely related to intent communication, which aims to allow observers to easily deduce the intent behind generated behaviors [14] [15]. Motions conveying intent have been shown to enhance perceived safety in virtual human-robot path-crossing tasks [16]. In reinforcement learning settings, robustness, efficiency and energy reward terms have been shown to enhance the ability of humans to interpret a robot’s intent [14]. Research focusing on projected visual legibility cues [17] has indicated that projected arrows generally offer greater interpretability than flashing lights in navigation settings. Notably, the means through which people convey and comprehend intent vary, and there exists no universally applicable method for assessing intent communication. By viewing communication as the exchange of information between agents, our TE-based framework holds the potential to adapt to various forms of communication, offering a versatile method to shape communication dynamics. Importantly, unlike many of the uni-directional intent communication approaches above, our framework captures the multi-directional nature of communication.
Transfer Entropy is a measure that allows the analysis of the information transfer and potential causal relationships between two simultaneous time series. TE is most prevalent in economics literature [18, 19], but has also been used to analyse animal-animal or animal-robot interactions  [20][21], joint attention [22] and to model pedestrian evacuation [23]. Berger et al. apply TE to detect human-to-robot perturbations using low-cost sensors [24]. TE has also been used to identify arbitrary cues from raw social interaction data between humans [25]. These works successfully quantify information transfer or the relationship between sources of information, but tend to focus on specific features or aspects of interest. Our work seeks to extend the use of TE to provide a more general approach aimed at enhancing information transfer in decision-making settings leading to improved human-robot collaboration.
Social learning: Mutual information (MI) is an information measure that captures the shared information between random variables, which has been widely used for social learning. Klyubin et al. [26] proposed the concept of empowerment for intrinsic motivation for reinforcement learning, which measures the maximum MI (channel capacity) between an agent’s actuators and their sensors. This concept was expanded by Mohamed and Rezende with a lower complexity maximisation approach to MI [27]. Jaques et al. use MI as a social influence reward for multi-agent deep reinforcement learning in Sequential Social Dilemmas (SSDs) [28] to encourage collaborations between agents [29]. While MI measures the shared information between random variables, TE measures the time-asymmetric information transfer. Unlike MI, the asymmetric nature of TE enables us to capture the directionality of information flow, which is advantageous for social communication. By leveraging TE, our goal is to quantify and modulate social information transfer.
3 Methodology
We consider an environment with multiple agents, each possessing distinct states, actions, and rewards. These rewards can be complementary (collaborative), competitive, or unrelated. The ego-agent aims to maximize its own rewards, ideally without negatively impacting other agents. In a HRI context, the robot should exhibit behaviour perceived as altruism. A fair outcome involves promoting collaboration, yielding to humans in competitive scenarios, and allowing independent agents to achieve their objectives if there is no conflict of interest. Within this environment, the ego-agent has complete access to its own state and can observe other agents, but may lack access to their decision-making processes or complete states. Consequently, we model interaction as a partially observable Markov decision process (POMDP), where states represent system situations and actions determine transitions between these states. Observations derived from states guide decision-making despite incomplete information.
The core concept behind the proposed method involves leveraging influence as an additional reward to enhance implicit communication between agents. This augmentation aims to accelerate learning and ultimately improve the collaboration performance for all agents or competitive capability of the other agents. Below, we assume two agents, labelled as $P1$ and $P2$ with $P1$ the ego-agent. By computing the TE from the historical actions of $P2$ (other agent) to the current action of $P1$ (ego-agent), P1 can better respond to P2 without an explicit model. This TE is computed as $TE_{\mathbf{a}_{2,t}^{(n)}\rightarrow a_{1,t}}$, where $a_{1,t}$ is ego-agent’s action at time $t$, $\mathbf{a}_{2,t}^{(n)}=(a_{2,t-1},a_{2,t-2},...,a_{2,t-n})$ comprises historical actions of $P2$, and $n$ is the history length. We augment the TE into the reward function,
$Reward_{a_{1,t}}=\phi TE_{\mathbf{a}_{2,t}^{(n)}\rightarrow a_{1,t}}+r.$ (1)
Here $r$ is the reward related to the agent’s objectives and $\phi$ is a scaling factor used to adjust the strength of the TE reward. Importantly, augmenting the reward with this TE implies encouraging the ego-agent to take actions that promote influence from past actions of the other agent to the current action of the ego-agent, making it more legible [15] to other agents. This occurs because of the ego-agent’s clearer response to the other agent’s actions, allowing us to affect other agents without requiring access to their behaviour models.
3.1 Measuring influence
We manipulate the influence using the Transfer Entropy, $T_{Y\rightarrow X}$, between the time series of state and action pairs. Transfer entropy is closely related to Wiener-Granger causality [30], and computes the conditional mutual information between two variables $X_{t}$ and $Y_{t}$ [31],
$T_{Y\rightarrow X}^{(k,l)}(t)$ $=I(X_{t}:\mathbf{Y}_{t}^{(l)}|\mathbf{X}_{t}^{(k)})=H(X_{t}|\mathbf{X}_{t}^{(k)})-H(X_{t}|\mathbf{X}_{t}^{(k)},\mathbf{Y}_{t}^{(l)})$ (2)
Here, $I$ denotes mutual information, while $H$ represents Shannon’s entropy [32]. This equation measures the directed transfer of information from $Y$ to $X$, with $t$ indicating time. The terms $\mathbf{Y}{t}^{(l)}$ and $\mathbf{X}{t}^{(k)}$ denote sequences of past observations for $Y$ and $X$, respectively, with lengths $k$ and $l$,
$\mathbf{X}_{t}^{(k)}=(x(t-\delta),x(t-2\delta),...,x(t-(k-1)\delta))$ (3)
$\mathbf{Y}_{t}^{(l)}=(y(t-\delta),y(t-2\delta),...,y(t-(l-1)\delta)).$
Here $\delta$ is the unit time step. In literature, $X$ is referred to as the ’target’ and $Y$ as the ’source’. Transfer entropy can be intuitively understood as the decrease in uncertainty within a state $X$, predicted solely based on its own past, upon the introduction of an additional information source $Y$ [31].
In order to compute TE, we calculate two entropy terms, the first assumes that there is influence from past actions of the other agent $P2$ to the current action of the ego-agent $P1$ (superscript +), while the second assumes no influence (superscript -). We denote $\mathbf{s}_{1,t}^{(n)}=(s_{1,t},s_{1,t-1},s_{1,t-2},...,s_{1,t-n})$ and $\mathbf{o}_{2,t}^{(n)}=(o_{2,t},o_{2,t-1},o_{2,t-2},...,o_{2,t-n})$, where $s_{1,t},s_{1,t-1},...,s_{1,t-n}$ and $o_{2,t},o_{2,t-1},...,o_{2,t-n}$ comprise the states of the ego-agent $P1$ and observations of agent $P2$ at time $t,t-1,...,t-n$ respectively. We then compute probability distributions over the actions (policies) for each of these scenarios:
$P_{a_{1,t}}^{+}$ $=P(a_{1,t}|\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)}),$ (4)
$P_{a_{1,t}}^{-}$ $=P(a_{1,t}|\mathbf{s}_{1,t}^{(n)})$
and calculate Shannon’s entropy. Here we show the formulation for discrete Shannon’s entropy, but Shannon’s differential entropy should be used if the action space is continuous. We can now calculate the TE, or information flowing from past actions of $P2$ to the current action of $P1$.
$TE_{\mathbf{a}_{2,t}^{(n)}\rightarrow a_{1,t}}$ $=TE_{f(\mathbf{o}_{2,t}^{(n)})\rightarrow a_{1,t}}=H(P_{a_{1,t}}^{-})-H(P_{a_{1,t}}^{+})$ (5)
$=[-\sum_{p\in P_{a_{1,t}}^{-}}p(a_{1,t})\log_{2}p(a_{1,t})]-[-\sum_{p\in P_{a_{1,t}}^{+}}p(a_{1,t})\log_{2}p(a_{1,t})].$
Our hypothesis is that manipulating TE will allow us to control the influence during social interactions. For instance, taking actions to increase TE could promote influence, which could sequentially enhance collaboration between agents. Ideally, being responsive and legible would aid all participants to achieve their goals during interactions. However, this strategy may become self-sacrificing or altruistic in some settings, a phenomena often explored in game theory [33]. This happens when being legible conflicts with an agent’s primary goal, or when an interaction is competitive.
Entropy measures the degree of uncertainty in actions: higher entropy means greater uncertainty, and lower entropy means less. Increasing $TE_{\mathbf{a}_{2,t}^{(n)}\rightarrow a_{1,t}}$ increases the uncertainty of actions when ignoring other agents and decreases the uncertainty when considering them (i.e., $H(P_{a_{1,t}}^{-})$ and $H(P_{a_{1,t}}^{+})$ in Equation (5) respectively), thus promoting influence. Conversely, decreasing TE resists this influence.
3.2 Q-learning
The TE calculations above require distributions over actions (a policy). We illustrate how these can be obtained using Q-learning [34]. However, it is important to highlight that our approach is not restricted to Q-learning and could be used with other probabilistic policies obtained through reinforcement learning. We employ Temporal Difference learning [35] to find Q-values representing the expected cumulative rewards for actions in specific states, following the Bellman equation [36],
$Q(s,a)\leftarrow(1-\alpha)Q(s,a)+\alpha(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})).$ (6)
Here $Q(s,a)$ is the Q-value for state $s$ and action $a$, $\alpha$ is the learning rate, $r$ is the immediate reward, $\gamma$ is the discount factor, $s^{\prime}$ is the next state and $a^{\prime}$ is the action in the next state. This iterative process enables the agent to learn the optimal Q-values over time, determining the most advantageous actions to take in various states. The optimal action $a^{*}$ at current state is determined by
$a^{*}=argmax_{a}Q(s,a).$ (7)
In contrast to this deterministic policy, a probabilistic policy can be employed by converting the Q values to probabilities, then sampling from the distribution to choose the action. We do this by taking the softmax [37] over Q-values at a given state.
$a^{*}\sim P_{a}=Softmax(Q(s,a)|_{s})$ (8)
Q-learning often uses an epsilon-greedy strategy [38] to balance exploration and exploitation. This approach alternates between choosing actions with the highest Q-value and occasionally exploring other options with a small probability ($\epsilon$). We use this strategy in our implementations.
In the proposed multi-agent setting where states are not fully observable, $s$ in Q-learning (referring to $Q(s,a)$) becomes the states of ego-agent and observations of other agents’ states in the scene. To be specific, $s$ comprises $\mathbf{s}_{1,t}^{(n)}$ and $\mathbf{o}_{2,t}^{(n)}$, which denote the states of $P1$ and observations of $P2$ at time $t,t-1,...,t-n$ respectively. The action $a$ is simply the actions of the agent $P1$ at time $t$ (i.e., $a_{1,t}$), resulting in the expanded Q table,
$Q(s,a)=Q(a_{1,t},\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)}),$ (9)
and probabilistic Q-learning policy
$a_{1,t}^{*}\sim P_{a_{1,t}}=Softmax(Q(a_{1,t}|\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)})).$ (10)
This distribution over actions allows us to compute TE using the Q table. First, for the entropy assuming influence, we can directly use the original Q table from Equation (10) as the probability distribution over actions:
$P_{a_{1,t}}^{+}=Softmax(Q(a_{1,t}|\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)})).$ (11)
In order to measure the entropy assuming no influence, the other agent’s ($P2$’s) history in the Q table can be marginalised over:
$Q(a_{1,t},\mathbf{s}_{1,t}^{(n)})=\frac{1}{m}\sum_{\mathbf{o}_{2_{t}}^{(n)}}Q(a_{1,t},\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)}),$ (12)
with $m$ the number of possible observations $\mathbf{o}_{2,t}^{(n)}$. Then the same process as above is repeated to compute the action distributions assuming no influence:
$P_{a_{1,t}}^{-}=Softmax(Q(a_{1,t}|\mathbf{s}_{1,t}^{(n)})).$ (13)
We can now calculate the TE according to Equation (5) and incorporate this into the reward term in the POMDP according to Equation (1).
4 Experiment and Results
We investigate the proposed approach in our corridor dilemma setting (Fig. 1), a simplified 11x5 grid-world where two players start at opposite ends of a corridor with their own randomly assigned objectives: passing or meeting. In each round, players move forward and choose to move left, straight, or right. The game ends when both agents reach the same row, resulting in either a pass or a meeting, with scores based on objectives. This game can be both competitive and collaborative, requiring anticipation, strategy, and potential collaboration.
Figure 1: A corridor dilemma episode: $P1$ is assigned to pass, and $P2$ is assigned to meet. In the first turn, $P1$ chose right, and $P2$ chose straight. After 5 turns, the result could be either pass (as in the top scenario), making $P1$ the winner, or meet (as in the bottom scenario), making $P2$ the winner.
We apply the proposed approach using agent coordinates as states and observations. In this specific application, the reward $r$ in Equation (1) is $r=0$ when the agent takes an action before reaching the middle row of the corridor, $r=10$ when the agent successfully achieves its objective, and $r=-10$ when it fails to achieve its objective after reaching the middle row. During the experiments, the TE reward components are normalised and scaled between 0 and 10 to match the objective rewards.
We conducted both simulation and human-agent experiments to validate the proposed method. For all experiments, we set both the Q-learning rate and discount factor to $0.8$. To avoid memory constraints and improve computational speed, we use sparse matrices for the Q tables in Eq. (12). A history length of 5 is used. ($i.e.,n=5$, the maximum history length.) Based on the TE reward scaling factor $\phi$, we define three types of agent: Non-TE agent ($\phi=0$, the baseline), Positive-TE agent ($\phi=10$, promoting influence) and Negative-TE ($\phi=-10$, resisting influence) agent.
4.1 Simulation
We conduct a range of simulations, putting each agent type against one another in a self-play framework where pairs of agents are simultaneously trained against each other for 30000 episodes with randomised objectives. $P1$ and $P2$ refer to the two players shown in Figure 1. For each experiment, we ran 30000 episodes and gathered results from six trials with distinct random seeds and compute their averages for presentation. More simulation details can be found in Appendix 6.1.
We report the averaged success rates for both competition and collaboration (denoted as SRCP and SRCL) to assess the agents’ capacity to compete or collaborate effectively. In this context, competition refers to scenarios where agents are assigned different objectives, while collaboration refers to situations where both agents share the same objective. These success rates are quantified as the number of successful outcomes in either competition or collaboration, divided by the total number of instances of competition or collaboration. The results are presented in Table 1.
Table 1: Final average success rate of competition (SRCP) and collaboration (SRCL) in simulation.
Experiment Agent (TE type) SRCP (%) SRCL (%)
Mean Std Mean Std
Random vs. Random $P1$ (Random) 50.00% - 50.00% -
$P2$ (Random) 50.00% -
Non-TE vs. Non-TE $P1$ (Non-TE) 48.31% 7.42% 63.50% 10.76%
$P2$ (Non-TE) 51.69% 7.42%
Non-TE vs. Pos-TE $P1$ (Non-TE) 63.38% 7.62% 76.97% 4.19%
$P2$ (Pos-TE) 36.62% 7.62%
Non-TE vs. Neg-TE $P1$ (Non-TE) 51.03% 8.16% 51.39% 5.90%
$P2$ (Neg-TE) 48.97% 8.16%
Pos-TE vs. Pos-TE $P1$ (Pos-TE) 49.95% 6.47% 91.72% 3.97%
$P2$ (Pos-TE) 50.05% 6.47%
Pos-TE vs. Neg-TE $P1$ (Pos-TE) 43.49% 7.96% 65.50% 10.62%
$P2$ (Neg-TE) 56.51% 7.96%
Neg-TE vs. Neg-TE $P1$ (Neg-TE) 49.84% 9.35% 49.70% 8.15%
$P2$ (Neg-TE) 50.16% 9.35%
In competition, the Non-TE agent in the Non-TE vs. Positive-TE experiment achieved the highest final performance with a 63.38% SRCP, with a corresponding reduction in SRCP for the paired Positive-TE agent. This is as expected, since increasing influence or legibility can lead to the self-sacrificing or altruistic behaviour described in section 3.1. This increase in influence allows the Non-TE agent to learn to satisfy its objectives more frequently both in circumstances of collaboration ($P1$ and $P2$ share the same objectives) and competition ($P1$ and $P2$ share different objectives).
Note that the Positive-TE vs. Positive-TE pair reaches the highest SRCL (91.72%), and also much fairer outcomes in terms of competition (49.95% vs. 50.05%) compared to the Non-TE vs. Positive-TE pair (63.38% vs. 36.62%). In addition, the Negative-TE agent in Positive-TE vs. Negative-TE group that tries to resist influence does not achieve a higher SRCP than the Non-TE agent in the Non-TE vs. Positive-TE group. These findings suggest that engaging in either receiving or resisting influence may compromise performance, but training with Positive-TE agent is beneficial for group collaboration and contributes to the fairness of the group. This is clearly evident as all pairs featuring at least one Positive-TE agent, consistently attain better collaboration performance than those without, and results in fairer competition outcomes (SRCP) for the Positive-TE vs. Positive-TE pair.
These experiments clearly show that performance in interactive settings can be affected by controlling influence in basic social interactions, indicating potential applicability to broader human-robot-interaction scenarios. Additional simulation results and analysis can be found in Appendix 6.2 and 6.3. Next, we verify the efficacy of the proposed method for agents interacting with human users.
4.2 Human-Agent Experiment
We conducted a user study in the corridor dilemma environment (See Appendix 6.4 for the interface design). Here, human participants played against three agents. Typically, individuals display ambivalence, occasionally leaning slightly towards selfishness or altruism, but not to an extreme degree. Therefore, a Non-TE agent, being ambivalent, could most accurately represent a typical human. Thus, participants are tested against three agents pre-trained under this assumption: 1. Non-TE agent ($P1$) vs. Non-TE agent ($P2$), 2. Non-TE agent ($P1$) vs. Positive-TE agent ($P2$), and 3. Non-TE agent ($P1$) vs. Negative-TE agent ($P2$). Each participant was asked to play against each of the three agents (randomly ordered) for 100 consecutive rounds with randomly generated objectives. The number of rounds was selected to provide ample opportunities for each participant to familiarise themselves with an agent and allow user performance to converge to a stable level. As part of the study, participants were told that one of the opponents could potentially be a real human player. We recruited 26 participants (11 female, 15 male, aged between 20 and 37). Their collaboration and competition performances were observed across 100 rounds of gameplay against each opponent, and subsequently averaged across all participants. These findings are illustrated in Figure 2.
Figure 2: User study results depicting collaboration and competition performance of humans. The first column shows average success rates, the second column displays violin plots of the final 20 rounds, where the mean values are marked, and the last column visualizes their distributions.
We conducted statistical analysis for the last 20 rounds, including Shapiro-Wilk tests [39] and Bartlett’s tests [40]. The data is normally distributed, but not homoscedastic. Therefore, we used Welch’s t-tests, which is reliable when the samples have unequal variances and possibly unequal sample sizes [41], and Games-Howell Post Hoc tests [42] to verify the significance of the findings. The outcomes from the Games-Howell tests indicate significant disparities in collaboration performance across all human players when engaging with the three distinct types of agents ($p\leq 0.05$). Regarding competition, significant differences were found in human performance when facing Positive-TE agents compared to encounters with Non-TE or Negative-TE agents. However, there was no significant difference ($p\geq 0.05$) in human performance when playing against Non-TE agents compared to interactions with Negative-TE agents.
It is clear that when playing against the Positive-TE agent, human participants achieve greater success rates in both collaboration and competition, when compared to other agents. This seems to be highly desirable behaviour in robots, agents should support humans to achieve their goals in collaboration, and behave transparently (and hence altruistically) when in competition. A Negative-TE agent resists influence, leading to decreased collaboration performance among human players. This results in competitive performance that remains relatively stable when compared to playing against Non-TE agents. Importantly, these real world findings match observations from the simulations conducted in Section 4.1. Figure 3 shows the performances of a trained Non-TE agents when facing the three different types of agents used in the human study. It should be noted that the simulation settings and Q-learning agents are considerably less complex than real human behaviour. Consequently, the effect of the TE reward in these simulations is notably more pronounced than in the user study, though both follow the same trends. Overall, these findings suggest that manipulating influence through TE does indeed impact human behavior. By boosting or resisting influence, we can design agents that facilitate improved or diminished performance in human participants.
Figure 3: Simulation results of Non-TE agents against three agent types. Rows depict collaboration and competition performance. Columns display average success rates, violin plots of the final 20 rounds, where the mean values are marked, and distributions of the last 20 rounds.
Figure 4: Partial survey results of the user study.
After each interaction, participants were asked to give feedback via a short survey about their experience with the agent. Figure 4 presents selected survey findings indicating participants perceive the Pos-TE agent as more legible and human-like. One-way ANOVA tests [43] showed no significant differences suggesting only subtle perceptual distinctions. Combined with the significantly different performance results in Figure 2, this implies that our method enables implicit influence modulation, or implicit communication. For more survey results, refer to Appendix 6.5. Additional findings, limitations and future work is discussed in Appendix 6.7.
5 Conclusion
This work proposed a framework to promote implicit communication for social HRI via influence modulation, which does not require explicit modelling or pre-existing knowledge of human participants. We test the proposed framework in both simulations and human-agent experiments. Results indicate that a positive or negative TE reward can enhance or reduce participant performance based on whether the interaction is collaborative or competitive. Notably, a positive TE reward fosters collaboration and altruism in competition, which is highly desirable for HRI. The framework requires only observations of other participants, and limited domain knowledge or assumptions to produce human-centric behaviours. As a result, this approach has significant potential for broader robotic applications.